X and z are independent. Var(x)=1 ; Var(z) = sigma ^2. E(z)= 0 and y= ax+b+z
I) cov(x,y)= ?
ii) corr(x,y)=?

dependent Varvane 2.
Answer #1

i) aV a

ii) Var(Y)= Var(aX+b+Z) = a^2Var(X) + Var(Z) = a^2 + sigma^2

and Corr(X, Y) =-/var(X)* Var(Y)-T*(a2 + σ*) al al